OpenStudy (anonymous):

Graph the line passing through the given point and having the given slope. Give the slope-intercept form of the equation of the line if possible. (2,-3),m=-3/5

hero (hero):

y = mx + b -3 = -3/5(2) + b -3 = -6/5 + b -3 + 6/5 = b -15/5 + 6/5 = b (-15 + 6)/5 = b (6 - 15)/5 = b -9/5 = b y = -3/5(x) - 9/5 y = (-3x - 9)/5